Bob and Steve want a 14% alcohol solution. how many ounces of a 12% alcohol solution must be mixed with a 20% alcohol solution to form 100 ounces of the 14% alcohol solution?

Answer by checkley79(3341)

.12X+.20(100-X)=.14*100
.12X+20-.20X=14
-.08X=14-20
-.08X=-6
X=-6/-.08
X=75 OZ. USED OF THE 12% SOLUTION.
100-75=25 OZ. OF THE 20% SOLUTION IS USED.
PROOF:
.12*75+.20*25=.14*100
9+5=14
14=14